### 17 Mar 2026 v2

# Guaranteeing Semantic and Performance Determinism in Flexible GPU Sharing

Abstract
GPU sharing is critical for maximizing hardware utilization in modern data centers. However, existing approaches present a stark trade-off: coarse-grained temporal multiplexing incurs severe tail-latency spikes for interactive services, while fine-grained spatial partitioning often necessitates invasive kernel modifications that compromise behavioral equivalence.
We present DetShare, a novel GPU sharing system that prioritizes determinism and transparency. DetShare ensures semantic determinism (unmodified kernels yield identical results) and performance determinism (predictable tail latency), all while maintaining complete transparency (zero code modification). DetShare introduces GPU coroutines, a new abstraction that decouples logical execution contexts from physical GPU resources. This decoupling enables flexible, fine-grained resource allocation via lightweight context migration.
Our evaluation demonstrates that DetShare improves training throughput by up to 79.2% compared to temporal sharing. In co-location scenarios, it outperforms state-of-the-art baselines, reducing P99 tail latency by 15.1% without compromising throughput. Furthermore, through workload-aware placement and our TPOT-First scheduling policy, DetShare decreases average inference latency by 69.1% and reduces Time-Per-Output-Token (TPOT) SLO violations by 21.2% relative to default policies.
**footnotetext: Equal contribution
1 Introduction
Artificial intelligence (AI) applications, such as chatbots [56, 69], code generation [40, 26], and autonomous agents [84], have become the dominant consumers of cloud compute resources and rely heavily on powerful GPU accelerators [20, 65, 47]. These workloads exhibit highly dynamic resource demands, frequently alternating between compute-bound phases (e.g., prompt prefill) and memory-bound phases (e.g., autoregressive decoding). Current cloud abstractions fail to match this dynamism, primarily offering rigid, fixed-size instances to enforce isolation. This mismatch compels operators to over-provision for peak loads, resulting in significant resource waste during off-peak cycles. For instance, Microsoft reports that compute utilization drops to under 10% during the memory-bound decoding phase when serving Llama3-8B [18] on A100 GPUs [43]. Industry surveys indicate that 60%–70% of GPU capacity routinely remains idle across both training and inference workloads in production clusters, translating to billions of dollars in wasted capital expenditure annually. Attempts to mitigate this inefficiency via manual resource tuning prove impractical, as the process is labor-intensive and fails to scale [39, 74].
GPU resource sharing offers a natural solution: by co-locating workloads, the capacity left idle during one task’s low-utilization phases can be reclaimed by co-resident jobs. Existing approaches can be broadly classified into two categories. Temporal sharing [75, 76, 77, 17] time-slices the GPU, granting each job exclusive access for fixed quanta. While straightforward, this approach fails to reduce under-utilization during memory-bound phases and can cause severe latency spikes for latency-critical services [22, 21]. Spatial sharing [15, 8, 30, 14, 83] permits concurrent execution by partitioning GPU resources (e.g., Streaming Multiprocessors (SMs)) across workloads. This approach is considered more promising for achieving high utilization, yet existing spatial-sharing systems remain impractical due to three main challenges:
  • C1: Semantic Determinism—Inequivalence after sharing. Fine-grained spatial sharing approaches [15, 9] often modify kernel code, fundamentally altering execution semantics. They change either the effective batch size (distorting normalization statistics) or the computation order (violating floating-point associativity), causing numerical divergence from standalone execution [32].
  • C2: Performance Determinsim—Unpredictable SLO violations. Co-locating services on shared GPUs inherently introduces resource contention. Compounding this, existing infrastructure [79, 58] suffers from loss of semantic information: it remains oblivious to request priority. This uniform handling exacerbates interference, jeopardizing strict tail-latency Service Level Objectives (SLOs).
  • C3: Transparency and Practicality—The engineering tax. Techniques like searching for fusible kernels [67] or manual kernel rewriting [9] require deep domain expertise and invasive code modifications, imposing an engineering tax. This not only creates maintenance burdens across rapidly evolving model architectures and underlying software stacks, but also amplifies the risk of violating semantic determinism (C1).
This paper presents DetShare, a transparent GPU sharing system that simultaneously addresses all three challenges. Our key observation is that the modern GPU software-hardware stack already offers the necessary primitives for efficient spatial sharing without sacrificing semantic determinism: NVIDIA’s Multi-Process Service (MPS) [10] enables spatial sharing with address space isolation, while Green Contexts [55] serve as lightweight execution units amenable to fine-grained preemption.
Leveraging these software-hardware mechanisms, DetShare introduces GPU coroutines, a novel abstraction that fully decouples logical execution contexts (visible to applications) from physical GPU resources (managed by the CUDA runtime). Applications operate on virtual contexts (vCtx) that are dynamically bound to physical contexts (pCtx) provisioned with varying SM quotas. A centralized, workload-aware scheduler orchestrates this binding, migrating contexts between physical resources to match instantaneous workload priorities.
Critically, by executing kernels without any modifications, DetShare ensures they yield identical results whether running in isolation or under sharing, thereby preserving both functional correctness and semantic determinism (C1). To meet tail-latency SLOs, DetShare incorporates workload-aware scheduling policies that prioritize latency-critical kernels while simultaneously maximizing aggregate throughput (C2). Finally, DetShare remains fully transparent: no manual tuning or kernel annotation is required (C3).
We evaluate DetShare on diverse workloads spanning DNN training, large language model (LLM) inference, and co-located training-inference scenarios. Our results show that DetShare improves throughput by up to 79.2% for DNN training workloads compared to temporal sharing. When co-locating DNN training and large-model inference tasks, DetShare can reduce the P99 tail latency of inference by up to 15.1% compared to the baseline while preserving training throughput. By introducing workload-aware job placement policies, DetShare reduces the average inference latency by up to 69.1% under real-world production traces. Furthermore, with our TPOT-First scheduling policy, the TPOT SLO violation rate is reduced by an additional 21.2% relative to the default policy.
Contributions.
We make the following contributions.
  • We systematically classify existing GPU sharing mechanisms and identify three fundamental challenges that limit their applicability and efficiency.
  • We propose GPU coroutines, a novel abstraction that fully decouples logical execution contexts from physical GPU resources, enabling flexible and fine-grained multiplexing with determinism.
  • We design and implement DetShare, a high-performance, general-purpose GPU sharing system that outperforms state-of-the-art solutions in both DNN training and LLM inference, while maintaining strict SLO compliance.
2 Background and Motivation
This section reviews AI model serving applications and existing mechanisms to run them on GPU.
2.1 AI Model Serving
AI models are now ubiquitous across modern applications, powering search, personalization, automation, diagnostics, and conversational agents. And model serving have become primary consumers of computing resources [20, 65, 47]. Serving AI models involves two broad phases, training produces a model and inference makes the model’s predictions available to various applications.
Training includes pre-training, referring to large-scale, often self-supervised training on massive data to build foundation models [13, 6, 46]. These pre-trained foundation models then serve as general-purpose backbones for downstream applications in natural language understanding, conversational agents, summarization, image representation learning, and multimodal reasoning. Post-training further adapts foundation models to domain-specific tasks through fine-tuning [35, 16], instruction tuning [73, 60, 49] or personalization [87, 82]. In addition, model compression [31, 34] and quantization [5] techniques are applied to reduce memory footprint and improve inference latency. Another kind of training is online or continual training [57], incrementally updating the deployed models using fresh or user-generated data.
Inference entails the execution of trained AI models to generate predictions for online or batch tasks. Within this domain, LLM serving has emerged as a dominant workload, underpinning a diverse array of critical applications ranging from conversational agents [56, 69] and neural code synthesis [40, 26] to Retrieval-Augmented Generation (RAG) pipelines [44, 19] and autonomous agentic systems [84]. Typical LLM inference can be conceptually divided into a prefill phase, where the input context is encoded and key-value caches are initialized, and a decode phase, where tokens are generated sequentially, often with beam search or sampling strategies. These phases exhibit distinct resource utilization patterns: prefill is primarily compute-bound, characterized by high arithmetic intensity due to parallel processing of input tokens. In contrast, decode is memory-bound, constrained by memory bandwidth due to the low arithmetic intensity of auto-regressive token generation and the overhead of KV cache loading [1, 86, 43].
2.2 GPU Sharing
An enduring challenge in systems and architecture, GPU sharing has attracted a broad spectrum of research. Existing solutions broadly fall into two paradigms: temporal and spatial sharing.
Temporal sharing [75, 76, 77, 17] allows one to switch from one application to another. This offers strong isolation, while context switching introduces heavy overhead and since only one context is activated, overall GPU utilization can be low when the workload is small [2, 4].
Spatial sharing [75, 76, 77, 17] executes multiple workloads concurrently across partitioned GPU resources, spanning from fine-grained software multiplexing like CUDA Streams to coarse-grained hardware isolation. While CUDA Streams [8, 48, 59, 71] enable intra-application parallelism by issuing operations into distinct hardware work queues, they operate within a single CUDA context. Although this eliminates context switching overhead, the resulting unified address space enforces fate sharing: a fault in one stream crashes the entire context, offering zero fault isolation.
To provide stronger multi-tenant isolation, vendors offer dedicated spatial partitioning mechanisms like NVIDIA’s Multi-Instance GPU (MIG) and Multi-Process Service (MPS). MIG [11] achieves strict hardware-level isolation by dividing a single physical GPU into fully isolated instances, each with dedicated compute and memory bandwidth. However, this isolation comes at a severe cost: MIG relies on heavyweight reconfiguration mechanisms that necessitate device resets, incurring latencies on the order of hundreds of milliseconds [67, 45]. Furthermore, its coarse-grained partitioning enforces fixed resource slice sizes, inevitably leading to significant resource fragmentation. In contrast to hardware-centric partitioning, MPS [10] operates at the software level, facilitating process-level spatial multiplexing via a client-server architecture. By multiplexing command streams from multiple independent clients into a single shared hardware context, MPS eliminates context-switching overheads while successfully maintaining separate virtual address spaces for each process. This design makes MPS particularly beneficial for multi-tenant inference workloads where clients can safely co-exist within a context, effectively boosting overall utilization by absorbing leftover parallelism without the rigid reconfiguration penalties of MIG.
2.3 The Sources of Semantic Non-Determinism
Hardware-Level Floating-Point Non-Associativity.
At the lowest level, the root cause of numerical divergence is the IEEE 754 standard for floating-point arithmetic [25]. Due to finite precision (especially in truncated formats like FP16 and BF16), floating-point addition is non-associative: $(A+B)+C\neq A+(B+C)$. Consequently, any variation in the sequence of arithmetic reductions will yield divergent rounding errors.
System-Level Serving Optimizations.
To maximize throughput, modern LLM serving engines dynamically adapt execution paths, which inherently introduces non-determinism. Techniques such as continuous batching [79] cause effective batch sizes to fluctuate heavily. Concurrently, high-performance computational kernels (e.g., cuBLAS) leverage dynamic tile sizing and Split-K reductions for matrix multiplications [32, 54]. By continually reshaping the underlying reduction trees, these performance-oriented optimizations alter the sequence of floating-point operations, directly exacerbating numerical divergence.
Application-Level Impact: Training-Inference Mismatch.
This numerical non-determinism extends beyond mere theoretical artifacts; it severely impacts advanced training paradigms such as Reinforcement Learning from Human Feedback (RLHF). During RL fine-tuning, identical prompts must consistently yield identical token trajectories. Micro-divergences introduced during inference trigger a policy shift [61]—the model’s behavior during online rollouts deviates from the policy evaluated during offline training. This mismatch effectively degrades online RL algorithms into suboptimal offline RL, substantially stalling convergence.
Recognizing this, recent serving frameworks (e.g., SGLang [85, 51]) and libraries [78] have introduced strict “deterministic modes”. However, these application-level guarantees implicitly assume dedicated, unfragmented hardware. As we discuss in Section 3.1, when these workloads are subjected to spatial multiplexing and sharing mechanisms at the infrastructure layer, the semantic determinism meticulously preserved by the application is once again destroyed.
3 Challenges of GPU Sharing Mechanisms
Table 1: Characterization of Existing GPU Sharing Mechanisms.
Sharing Mechanism Key Techniques Semantic Determinism Performance Determinism Transparency & Practicality Utilization
Temporal Sharing Time-slicing Low
Spatial Sharing Kernel atomization or scheduling High
Hybrid Sharing Combination of both Medium
  • Note: ◗ Partially addressed, ❍ Fully addressed, ⚫ Not addressed.
In this section, we classify and summarize these existing GPU sharing mechanisms, as shown in Table 1.
Current GPU sharing mechanisms can be broadly categorized into temporal sharing, spatial sharing, and hybrid sharing approaches. While spatial sharing methods generally outperform temporal sharing in terms of utilization and SLO compliance, they often require significant human effort to tune and may not guarantee equivalence after sharing.
3.1 Challenge-#1: Semantic Determinism
The bedrock of production AI systems is semantic integrity: a model validated in an offline environment must behave identically when deployed online. The core requirement for GPU resource sharing, therefore, is behavioral equivalence—preserving each workload’s functional behavior during shared execution compared to its standalone counterpart. This property is paramount for training-inference consistency [50]: models validated on dedicated hardware must yield identical predictions in inference scenarios. Any numerical divergence—however minute—can invalidate offline accuracy metrics. While temporal sharing naturally preserves arithmetic determinism by granting exclusive access to full resources, spatial sharing introduces complex challenges.
To maximize utilization, recent intrusive methods (e.g., LithOS [9]) employ kernel atomization—dynamically partitioning parallelizable dimensions to fit available resource fragments. We argue that such techniques fundamentally violate the abstraction boundary between logical computation and physical resource availability. Specifically, standard GPU kernels operate on a contract where the result is a function of input data ($D$) and model state ($S$): $Result=F(D,S)$. However, by dynamically reshaping kernels to fit hardware fragments, atomization implicitly leaks instantaneous resource constraints ($R(t)$) into execution semantics, effectively altering the function to $Result^{\prime}=F(D,S,R(t))$. Since $R(t)$ is stochastic in a multi-tenant cluster, the computation becomes non-deterministic. We categorize this divergence into two distinct levels:
Figure 1: Statistical Divergence. The impact of dynamic batch fragmentation on training loss. Adapting batch sizes to fit resource fragments alters the statistical properties of the workload, leading to increased loss variance compared to the baseline.
  • Statistical Divergence (Batch Sensitivity): Techniques that partition resources by splitting batches effectively alter the statistical population used for operations like Batch Normalization [38]. As demonstrated in Figure 1, reducing the effective batch size to fit fragmented resources distorts mean and variance statistics, leading to significant loss fluctuations and degraded convergence.
  • Numerical Divergence (FP Non-associativity): A more insidious issue arises in reduction-heavy kernels (e.g., Softmax, LayerNorm). Due to the non-associativity of floating-point arithmetic [25], the order of summation determines rounding error accumulation. As shown in Figure 2, we simulate resource fragmentation by splitting a large tensor summation into varying grid sizes. The results demonstrate that merely altering the reduction tree structure causes the absolute error (vs. standalone execution) to fluctuate significantly. This divergence is present in both FP16 and BF16, with BF16 showing an order of magnitude higher error, confirming that fragmentation breaks bitwise determinism.
Figure 2: Numerical Divergence. Absolute numerical deviation of FP16 and BF16 summations compared to their unfragmented, standalone execution. Dynamically splitting the computation grid (X-axis) inevitably alters the underlying reduction tree, exposing the non-associativity of floating-point arithmetic. This semantic divergence is especially severe for BF16.
Formalization.
We formalize this observation as the Resource-Coupling Risk. Let $\Phi$ be a deterministic kernel function. For any two execution instances $i$ and $j$ with identical inputs $x$, the condition $\Phi(x)_{i}\equiv\Phi(x)_{j}$ must hold. However, intrusive partitioning introduces a resource mapping function $M:\mathcal{K}\rightarrow\mathcal{R}_{avail}$, where the kernel implementation $K$ adapts its reduction strategy based on $M$. Consequently, the output divergence $\delta$ becomes non-zero:
$\delta=|K(x,M(R_{i}))-K(x,M(R_{j}))|>0$ (1)
In deep pipelines like LLM inference, these accumulated "micro-divergences" create a validation-deployment gap unacceptable for production critical paths.
Principle 1: Kernel execution semantics must remain unchanged to guarantee sharing equivalence.
3.2 Challenge-#2: Performance Determinism
Meeting strict tail-latency SLOs is a prerequisite for ensuring satisfactory user experience in interactive deep learning workloads. However, achieving performance determinism—specifically, keeping P99 latency stable—is notoriously difficult in multi-tenant environments. While temporal sharing is the incumbent solution, it is fundamentally ill-suited for strict latency constraints. The primary limitation is that time-slicing enforces serialization: it introduces unavoidable queuing delays for applications not currently scheduled on the GPU [27, 77]. As contention increases, these accumulated wait times frequently cause tail latencies to exceed strict deadlines.
Theoretical Containment ($\mathcal{S}_{temporal}\subset\mathcal{S}_{spatial}$).
Spatial sharing offers a fundamentally more promising approach. We observe that any performance target achievable via temporal sharing is theoretically achievable via spatial sharing. Formally, let $\mathcal{S}_{temporal}$ be the set of valid schedules achievable via time-slicing (exclusive ownership for $\Delta t$) and $\mathcal{S}_{spatial}$ be the set of spatially partitioned schedules. Since temporal sharing is mathematically a degenerate case of spatial sharing—where partition sizes toggle strictly between $0\%$ and $100\%$—it follows that $\mathcal{S}_{temporal}\subset\mathcal{S}_{spatial}$.
Corollary: If a workload set can satisfy SLOs under a temporal policy, there theoretically exists a spatial configuration that also satisfies these SLOs.
The Reality Gap: Unmanaged Interference.
Despite this theoretical potential, translating spatial partitioning into practical service assurance is hindered by performance non-determinism, which we categorize into two distinct domains:
  • Micro-architectural Contention: Spatial co-location implicitly multiplexes shared resources that are not explicitly partitioned, most notably Device Memory Bandwidth and Last-Level Cache capacity. As characterized in prior studies [22], a latency-sensitive inference job can suffer non-deterministic latency spikes (up to $3\times$ [43]) when co-located with a bandwidth-heavy training job.
  • Scheduler Obliviousness: Commodity GPU hardware schedulers are throughput-oriented and oblivious to high-level SLO constraints. Without explicit intervention, a high-priority kernel may suffer from head-of-line blocking behind a massive, non-preemptible background kernel in the hardware command queue [64], rendering logical priority ineffective.
While spatial sharing theoretically encompasses the solution space for meeting SLOs, naive partitioning fails to guarantee performance determinism due to unmanaged resource contention. To bridge this gap, the sharing mechanism requires an SLO-aware scheduling policy coupled with low-latency kernel preemption, ensuring that high-priority queries can immediately reclaim resources from best-effort workloads.
Principle 2: The scheduling framework should be SLO-aware and support low-latency preemption.
3.3 Challenge-#3: Transparency and Practicality
We argue that widespread adoption of GPU sharing hinges on deployment transparency: the infrastructure must optimize resource usage without imposing any semantic or syntactic constraints on the application layer.
Ideally, the resource management runtime should strictly adhere to the end-to-end principle, treating Deep Learning models as opaque executables (i.e., "Black Boxes"). This decoupling allows for "Day-Zero" support: new model architectures can be deployed immediately without engineering intervention.
However, many existing spatial sharing designs [15, 67] sacrifice this practicality for theoretical utilization gains. Approaches relying on static compiler analysis, operator fusion, or manual kernel rewriting introduce a high "Engineering Tax". By requiring deep introspection into model graphs or invasive source code modifications, these methods tightly couple the sharing substrate with specific model implementations. This coupling creates a brittle ecosystem where even minor updates to framework or model architecture can invalidate the optimization logic, necessitating expensive manual retuning.
The Hidden Cost: Verification Explosion.
Crucially, this coupling imposes a burden beyond development—it fundamentally expands the verification surface. In hyperscale production environments, maintaining offline-to-online equivalence is non-negotiable. Mechanisms that modify user GPU kernels or the runtime environment effectively invalidate rigorous offline validation processes. This operational reality is corroborated by infrastructure practitioners at ByteDance [7], who note that "mechanisms requiring code rewriting are typically rejected not just for complexity, but because they mandate a prohibitive re-verification process for every framework or model update."
Principle 3: The sharing mechanism must be transparent, precluding any modification to user code or model semantics.
4 DetShare Design
This section details the design of DetShare, a system that reconciles execution determinism with high hardware utilization. Existing spatial sharing approaches often compromise semantic determinism by transparently altering block size to fit fragmented resources [9]. In contrast, DetShare adopts a transparent spatial multiplexing approach. The core design philosophy of DetShare is the strict decoupling of logical execution flows from physical resource provisioning. By maintaining kernel launch configurations as immutable invariants, DetShare ensures that resource elasticity never alters the arithmetic behavior of the application.
We first present the overall system architecture and the core GPU Coroutine abstraction (§4.1). We then detail the runtime mechanisms for dynamic context binding and cooperative preemption (§4.2), followed by the state management optimizations that enable low-overhead migration (§4.3). Finally, we describe the extensible scheduling framework (§4.4) and our tiered fault isolation strategy (§4.5).
4.1 Overview
Figure 3: The overall architecture of DetShare.
As illustrated in Figure 3, DetShare introduces a dynamic decoupling layer between the application and the GPU hardware. The core abstraction enabling this flexibility is the GPU Coroutine, which decouples the application’s logical view of the device from the underlying physical hardware constraints. DetShare formalizes this decoupling through two primary entities:
  • Virtual Context (vCtx): A persistent, application-facing handle that encapsulates the logical execution state. DetShare exposes vCtx as a fully transparent abstraction, functionally equivalent to a standard CUDA context from the client’s perspective.
  • Physical Context (pCtx): An ephemeral, hardware-backed resource container. Each pCtx is a strictly isolated execution unit provisioned with a hard limit on SMs.
Determinism via Abstraction. By separating the logical definition of a task from its physical execution capacity, DetShare guarantees semantic determinism. When resource availability fluctuates, DetShare does not perform kernel atomization (which risks numerical divergence). Instead, it employs a migrate-to-yield mechanism: the vCtx is transparently rebound to a different pCtx with an adjusted resource quota. Consequently, the computational function remains $Result=F(D,S)$, independent of the instantaneous resource state $R(t)$.
4.2 Mechanism: GPU Coroutines
The DetShare runtime manages the lifecycle of GPU Coroutines, handling dynamic binding, execution signaling, and cooperative preemption.
The pCtx Pool.
Dynamically instantiating hardware contexts incurs significant driver latency, which is prohibitive for microsecond-scale scheduling. DetShare mitigates this by maintaining a pre-provisioned pool of pCtx instances at initialization. These contexts are configured with discrete SM quota tiers (e.g., 25%, 50%, 100%). This pooling strategy removes control-plane overhead from the critical path, enabling zero-allocation-latency dispatch and migration.
Dynamic Binding and Dispatch.
DetShare maintains a Binding Table to track the mapping between active vCtxs and physical resources. Upon intercepting a launch API (e.g., cuLaunchKernel), the runtime consults the global scheduler to determine the optimal placement. Depending on the current mapping and resource state, the dispatch process proceeds under two distinct scenarios:
  • Direct Dispatch: If the vCtx is already bound to a pCtx that meets current SLO requirements, the kernel is directly enqueued to the hardware command queue. This execution flow incurs negligible overhead.
  • Context Remapping: If the scheduler detects a priority inversion or resource fragmentation, it triggers a Remapping Event prior to execution. The vCtx is unbound from its current container and undergoes a lightweight migration to a target pCtx with the appropriate quota.
Cooperative Preemption via Signaling.
To support latency-sensitive workloads, DetShare implements a low-latency preemption mechanism that strictly preserves kernel semantics. As shown in Figure 4, DetShare injects a lightweight Resident Control Kernel (RCK) into each active pCtx. The RCK is a persistent, single-thread kernel consuming negligible resources ($0.1\%$ SM occupancy), acting as a device-side signal handler.
When the scheduler necessitates preemption (e.g., to reclaim SMs for a high-priority inference task), it asserts a flag in a shared memory region mapped to the target pCtx. The RCK detects this signal and triggers a cooperative yield. Unlike traditional mechanisms that terminate the executing kernel, DetShare leverages CUDA Stream dependencies to pause execution at the nearest kernel or instruction block boundary. This safe suspension allows DetShare to synchronize state and initiate the context migration process, effectively yielding physical resources while preserving the logical progress of the preempted task.
Figure 4: An example of Remapping Event and pCtx migration.
4.3 State Management and Migration
Minimizing the volume of data transferred between pCtx instances is critical for efficient context migration. Since inter-context memory movement typically dominates migration latency, DetShare employs three complementary optimizations to mitigate this overhead: fine-grained working set tracking, dependency-aware input analysis, and demand-driven lazy copying.
Fine-Grained Usage Tracing.
A naive migration strategy that snapshots the entire GPU memory address space incurs prohibitive overheads, particularly for GPU kernels that access only a fraction of their allocated buffers. To address this, DetShare implements fine-grained tracing to monitor the "dirty" state and access patterns of memory regions. By maintaining a precise view of the active memory footprint for each vCtx, the runtime identifies the exact subset of data requiring coherence in the destination pCtx, thereby filtering out stale or idle pages from the migration set.
Introspection-Based Input Analysis.
To further prune the migration footprint, DetShare performs lightweight data-flow analysis on GPU kernel launch parameters. Prior to dispatch, the runtime intercepts execution to examine parameter structures and their associated device pointers, deriving the minimal closure of memory objects required for correctness. By isolating the kernel-reachable working set, DetShare avoids the indiscriminate transfer of the global heap, restricting data movement strictly to the regions touched during execution. This approach significantly reduces migration latency, especially for workloads exhibiting sparse access patterns over large memory allocations.
Demand-Driven Lazy Copying.
Finally, DetShare leverages a lazy migration policy to amortize data transfer costs. During a migration event, only the critical execution state–comprising GPU kernel launch parameters and their immediate dependencies—is eagerly transferred to the target pCtx. The remainder of the working set is migrated on demand: if the resumed kernel accesses a non-resident region, DetShare triggers an asynchronous, incremental transfer in the background. This strategy eliminates synchronous blocking, effectively overlaps communication with computation, and ensures that migration remains transparent to the application’s control flow.
4.4 Extensible Scheduling Framework
DetShare does not enforce a monolithic scheduling algorithm. Instead, it adopts a decoupled architecture that strictly separates the scheduling mechanism (context switching, signal injection) from the scheduling policy. This design exposes an extensible framework, allowing operators to define custom logic tailored to specific application SLOs without modifying the underlying runtime.
The Pluggable Policy Interface.
DetShare abstracts the scheduling decision process into a Policy Interface. Users or cluster administrators can implement this interface to define how DetShare reacts to lifecycle events. The framework exposes key hooks such as OnLaunch, OnCompletion, and OnCongestion. By implementing these methods, custom policies can inspect the global state—including the active pCtx pool status, queuing delays, and historical kernel execution profiles—to make granular decisions.
For example, a user can implement a "Cost-Aware" policy that prioritizes throughput during off-peak hours but switches to strict latency prioritization when a specific flag is set. DetShare treats these policies as pluggable modules, enabling the runtime to switch strategies dynamically without recompilation.
Reference Policies.
While the framework is extensible, DetShare provides two production-grade reference policies:
  • SLO-Aware Policy (Default): This policy targets mixed inference-training workloads. It utilizes a lightweight predictor (based on historical grid sizes and duration) to estimate the "Head-of-Line Blocking" time. If an incoming high-priority kernel is predicted to miss its deadline due to a running background task, the policy returns a preempt decision, triggering the mechanism described in §4.2.
  • TPOT-First Policy: Designed specifically for generative AI, this policy (evaluated in §6.2) introspects the semantic phase of LLM inference. It prioritizes the decoding phase over the prefill phase, ensuring deterministic inter-token latency even under heavy contention.
QoS Enforcement Mechanisms.
Regardless of the chosen policy, DetShare operationalizes scheduling decisions through a robust QoS layer. This layer acts as a translation barrier, mapping high-level logical directives into concrete physical resource partitions. By leveraging hardware-assisted isolation primitives for both spatial partitioning and temporal throttling, DetShare ensures that the user-defined policies are strictly bounded by hardware-enforced guarantees, preventing low-priority workloads from eroding the performance isolation of critical tasks.
4.5 Fault Isolation
A robust multi-tenant system must ensure that failures in one context do not propagate to others. DetShare implements a two-tiered fault isolation strategy to handle both fatal exceptions (which compromise correctness) and soft hangs (which degrade performance).
Fatal Exception Containment.
Fatal exceptions are categorized based on their blast radius. Local exceptions (e.g., illegal memory access within a kernel) corrupt only a single pCtx. DetShare leverages MPS-based static partitioning to strictly contain these failures; a crash in one pCtx terminates only the bound vCtx, while kernels in concurrent pCtx instances continue execution uninterrupted. Global exceptions (e.g., uncorrectable ECC errors or MMU faults) may contaminate the entire GPU device state. To mitigate cross-context impact, DetShare employs an emergency migration protocol. Upon detecting a global fault signal, the runtime pauses all unaffected vCtx instances and seamlessly migrates them to standby pCtx slots on alternative GPUs. This cross-device migration ensures high availability even in the presence of catastrophic hardware failures.
Soft Hang Mitigation.
Soft hangs, such as deadlocks or infinite loops, represent non-fatal but resource-wasting stalls. The DetShare scheduler detects these conditions using timeout-based heuristics: if a kernel exceeds its predicted execution window by a configurable threshold (e.g., $3\times$), it is flagged as potentially hung. To prevent these pathological workloads from monopolizing resources, the scheduling policy progressively demotes the offending vCtx. Upon subsequent scheduling rounds, the hung kernel is preempted and migrated to pCtx instances with strictly capped SM quotas (e.g., minimal tier). This "quarantine" approach ensures that while the hung task is not immediately killed (preserving debugging state), its resource consumption is throttled to prevent performance degradation for well-behaved concurrent tasks.
5 Implementation
The core of DetShare comprises approximately 6,000 lines of C++ code and 1,000 lines of CUDA code. To ensure both application transparency and fault tolerance, the implementation is architected around two primary components: a lightweight interception shim layer and a standalone global scheduler daemon.
To interpose on workload submissions without modifying user codes, the shim layer (see Figure 3) is transparently injected via LD_PRELOAD to intercept CUDA Driver and Runtime APIs (e.g., cuLaunchKernel, cudaMalloc). These API wrappers capture launch metadata and seamlessly enqueue requests into lock-free, shared-memory ring buffers. By leveraging shared memory for Inter-Process Communication (IPC) with the global scheduler, DetShare minimizes submission latency and keeps the direct dispatch overhead negligible.
At the execution level, DetShare employs a hardware-software co-design to enforce strict physical resource isolation. The global scheduler instantiates physical GPU contexts (the pCtx pool) utilizing NVIDIA Green Contexts, which achieve fine-grained spatial partitioning by enforcing hard quotas on SMs. To safely multiplex these hardware units while strictly preserving the execution semantics of GPU kernels, DetShare orchestrates the contexts over NVIDIA’s MPS to guarantee robust address space isolation.
6 Evaluation
Our key takeaways from the evaluation are:
  • Higher Training Throughput: By enabling efficient spatial sharing, DetShare improves training throughput by up to 79.2% compared to temporal sharing.
  • Lower Inference Latency: With workload-aware job placement policies and lightweight context migration, DetShare reduces the average inference latency by up to 69.1%.
  • Flexible Scheduling Policy: Utilizing our pluggable TPOT-FIRST policy, DetShare reduces the TPOT SLO violation rate by an additional 21.2% relative to the default policy, demonstrating robust adaptability to dynamic traffic.
Experimental Setup.
We evaluate DetShare across two distinct hardware platforms to capture diverse resource bottlenecks: an NVIDIA A800 cluster and NVIDIA Hopper architecture GPUs (e.g., H20). To accurately reflect multi-tenant production environments, our workloads span a broad spectrum of AI tasks. Rather than executing tasks in isolation, we construct rigorous colocation scenarios by pairing diverse throughput-oriented training jobs together, and by co-scheduling heavy background training tasks alongside latency-critical LLM inference streams.
6.1 End-to-End Performance Evaluation
We evaluate the end-to-end performance of DetShare to demonstrate its efficacy in maximizing aggregate throughput while maintaining strict performance isolation. Our evaluation covers two distinct colocation scenarios: (1) Homogeneous and Heterogeneous Model Training (Table 2); and (2) Mixed Workload Colocation, pairing throughput-oriented training with latency-critical LLM inference (Table 3).
Baselines.
We compare DetShare against industry-standard hardware partitioning mechanisms (NVIDIA MPS and MIG), as well as state-of-the-art software-defined schedulers including Orion [67], Tick-Tock [71], and Salus [81]. Since the source code for LithOS [9] is not publicly available, we implemented a faithful reproduction of its scheduling logic based on its design principles to serve as a comparative baseline.
6.1.1 Colocated AI Model Training
Table 2: Experimental setup for Colocated AI Model Training. Workloads are categorized by their dominant resource bottleneck.
Category Config Colocated Workloads
Job 1 Job 2
High Compute Contention A ResNet-50 [33]
B ResNet-101
C BERT-Large [13]
D BERT-Large ResNet-50
Low Compute Contention 1 RNN [62]
2 DeepFM [29] ResNet-101
LLM Training E GLM-4-9B [23] (LoRA [36])
Figure 5: Normalized throughput of colocated model training tasks (higher is better). DetShare consistently outperforms baselines, particularly in scenarios with high resource contention or varying interference patterns.
Figure 5 reports the normalized throughput across all configurations. DetShare consistently achieves superior aggregate throughput by effectively balancing SM utilization and minimizing inter-job interference.
High Compute Contention (Configs A–D).
These configurations represent arithmetic-bound scenarios involving compute-intensive models where the GPU’s SMs are saturated. Traditional temporal sharing and coarse-grained spatial partitioning often degrade performance due to resource fragmentation. Taking Config C (dual BERT-Large) as an example, DetShare attains a normalized throughput of $0.58\times$ per job. In contrast, rigid partitioning schemes like MIG are limited to $0.53\times$, while temporal sharing (Orion) drops to $0.50\times$ due to high context-switching overheads. By leveraging pCtx pooling and dynamic vCtx-pCtx dispatch without stalling the pipeline, DetShare significantly reduces inter-job interference, maintaining high efficiency even under heavy arithmetic contention and outperforming static heuristics such as GSlice.
Low Compute Contention (Configs 1–2).
Workloads such as RNNs and DeepFM exhibit intermittent GPU utilization, often bounded by memory bandwidth or host-device I/O. DetShare effectively harvests these ephemeral idle cycles (“execution bubbles”) by rapidly migrating vCtxs to underutilized pCtxs. In Config 1 (dual RNN jobs), DetShare achieves near-optimal isolation with normalized speeds of $0.99\times$ per job, effectively matching dedicated execution performance. While baselines like Tick-Tock and Salus plateau around $0.85\times$$0.90\times$, DetShare exhibits superior scheduling granularity in identifying and filling millisecond-level idle slots.
LLM Training (Config E).
We further evaluate DetShare on large-scale foundation model fine-tuning (GLM-4-9B with LoRA). DetShare ($0.95\times$) outperforms hardware-level partitioning (MIG, $0.93\times$) by enabling flexible resource redistribution rather than static allocation. Notably, static spatial schedulers like GSlice struggle in this scenario ($0.75\times$$0.79\times$), failing to manage the massive kernel launches and dynamic memory footprints associated with LLMs. DetShare sustains high execution efficiency even under these intense resource demands, highlighting its robust applicability to modern GenAI workloads.
Table 3: Experimental setup for Mixed Workload Colocation. We pair a throughput-oriented batch job (Job 1) with a latency-critical inference job (Job 2).
Category Config Colocated Workloads
Job 1 (Train) Job 2 (Inference)
CV + LLM Inference A ResNet Llama 3 [18]
B ResNet GPT-J [70]
NLP + LLM Inference 1 BERT Llama 3
2 BERT GPT-J
6.1.2 DNN Training with LLM Inference
To evaluate DetShare in mixed-criticality scenarios, we pair throughput-oriented training jobs with latency-sensitive LLM inference requests. The specific workload combinations (Table 3) cover Computer Vision (CV) and NLP models as background tasks, colocated with Llama 3 and GPT-J inference streams.
Figure 6: Performance comparison on mixed workloads. (a) DNN training job throughput (higher is better) and (b) LLM inference job p99 latency (lower is better) under four configurations. DetShare effectively balances system efficiency and SLO, outperforming methods like Orion and LithOS in both metrics.
Figure 6 presents the performance trade-offs between batch training throughput and inference tail latency. DetShare demonstrates a superior Pareto frontier compared to all baselines, successfully maintaining strict SLOs for interactive workloads while maximizing the utilization of residual GPU cycles for background training.
Guaranteeing Inference SLOs.
For latency-critical LLM workloads, DetShare achieves the lowest 99th percentile (p99) latency across all configurations. In Config A, DetShare reduces the p99 latency to 4.5 (normalized units), representing a $2.6\times$ and $2.2\times$ reduction compared to Orion (12.0) and MIG (10.0), respectively. Similarly, in the more memory-intensive Config B (ResNet + GPT-J), DetShare maintains a p99 latency of 5.9, significantly outperforming Orion (15.0) and MPS (10.2). Hardware partitioning methods like MIG enforce inflexible spatial boundaries; they lack the dynamic prioritization required to accommodate bursty inference traffic. Conversely, existing software schedulers like Orion fail to provide strict performance isolation. Under coarse-grained task co-location, critical inference GPU kernels are easily starved by heavy training workloads, resulting in severe interference. DetShare, by contrast, couples compute resource isolation via Green Contexts with cooperative preemption driven by injected RCKs. This joint mechanism ensures strict performance isolation and immediate physical resource reclamation for latency-sensitive inference tasks, without violating semantic determinism.
Maximizing Training Throughput.
While prioritizing inference, DetShare does not compromise the progress of background training jobs (Job 1). For example, in Config 1, DetShare achieves a normalized throughput of 0.60, outperforming LithOS (0.58), MPS (0.55), and Orion (0.53). Because inference workloads exhibit bursty arrival patterns, static partitioning (MIG/MPS) leaves compute units underutilized during idle gaps. DetShare’s dynamic vCtx-pCtx remapping allows the background training job to instantly reclaim these idle cycles, maximizing aggregate GPU utilization.
Figure 7: End-to-end evaluation on Azure, LongBench, and BurstGPT traces: Throughput (higher is better), Latency Breakdown and SLO Violations (lower is better). DetShare reduces average and tail latency while maintaining throughput. Adopting the TPOT-First strategy further decreases decoding tail latency.
6.2 Scheduling Policy Evaluation
We evaluate the impact of DetShare’s extensible scheduling policies through a comprehensive case study.
6.2.1 Traces and Baselines
To demonstrate DetShare’s transparency and effectiveness at the infrastructure layer, we evaluate an unmodified instance of the state-of-the-art serving system, SGLang [85], running entirely on top of DetShare to serve the Llama-3.1-8B-Instruct model. We compare this against a baseline of SGLang executing on a default, exclusively provisioned GPU.
We run across three traces representing distinct production scenarios:
  • Azure LLM Trace 2024 [66]: Highly stochastic request arrival times, testing general adaptability.
  • LongBench [3]: Long-context tasks imposing significant pressure on GPU memory capacity and bandwidth.
  • BurstGPT [72]: Extreme traffic spikes challenging the scheduler’s queue management.
We analyze DetShare under two pluggable configurations: the Default Policy (balancing resource fairness) and the TPOT-First Policy (explicitly prioritizing the decoding phase).
6.2.2 Performance Analysis
Figure 7 depicts the throughput, latency breakdown, and SLO violation rates. Overall, DetShare-backed SGLang significantly outperforms the baseline. For instance, on the Azure trace, DetShare reduces the average TTFT by $69.1\%$ ($0.55$s to $0.17$s) while maintaining comparable throughput. Furthermore, DetShare demonstrates superior stability by reducing the P99 TPOT on LongBench by $4.7\times$ (1723.91723.9ms vs. 366.2366.2ms). Given that strict tail-latency guarantees during token generation are paramount for a seamless interactive user experience, this dramatic reduction highlights DetShare’s efficacy even in severely memory-constrained scenarios.
Impact of TPOT-First Policy.
The TPOT-First Policy exhibits enhanced robustness in minimizing generation latency violations, particularly under fluctuating loads.
  • General Improvement: On the Azure trace, the TPOT-First policy reduces TPOT SLO violations by $21.2\%$ compared to the Default policy.
  • Trade-off in Bursty Scenarios: The advantages of the TPOT-First policy are most pronounced in the BurstGPT benchmark. As shown in Figure 7, this policy reduces TPOT SLO violations by $46.1\%$ ($17.77\%$ to $9.58\%$) and dramatically lowers the P90 TPOT from 847.7847.7ms to 262.9262.9ms. However, we observe a concomitant increase in TTFT SLO violations, which rise from $4.33\%$ to $6.37\%$. Analysis: We attribute this phenomenon to a necessary scheduling trade-off. Under extreme burstiness, the TPOT-First policy proactively delays the scheduling of new prefill requests (increasing TTFT) to reserve compute and memory bandwidth for ongoing decoding phases. This strategy effectively insulates active requests from interference caused by sudden spikes in arrival rates, ensuring a smooth generation experience at the cost of slightly higher initial queuing delays.
Figure 8: Overhead analysis. DetShare incurs only 4% and 12% overhead for context switching and preemption, respectively, normalized to the exclusive execution.
6.3 Overhead Analysis
Figure 8 quantifies the performance overhead introduced by DetShare, normalized to the exclusive GPU baseline. Overall, DetShare incurs a marginal overhead of 4% for context switching and 12% for preemption.
Analysis.
The context switching overhead is kept minimal because the pCtx leverages Context Memory Tracing to avoid redundant data migration between device and host memory. The 12% preemption overhead is primarily dominated by the internal synchronization costs when the RCK interrupts the GPU execution flow. Given the substantial reductions in tail latency and SLO violations (e.g., $4.7\times$ lower P99 TPOT), we argue that this moderate overhead is a highly favorable trade-off for multi-tenant production environments.
7 Related Work
Temporal Sharing.
Temporal sharing remains the default mechanism for multi-tenant GPU execution. Prior systems have significantly enhanced this approach for DNN workloads by mitigating context-switching overheads through state reuse (Salus [81]), dynamically managing containerized or runtime memory footprints (Antman [77], TGS [75], Zico [48]), and employing predictive scheduling to bound inference tail latency (Clockwork [27], Gandiva [76]). However, these solutions inherently operate at the granularity of the entire GPU. Consequently, they enforce strict execution serialization and fail to overlap computation. This limitation prevents them from reclaiming idle capacity during memory-bound phases or bursty workloads, leaving significant hardware resources underutilized.
Spatial Sharing.
Spatial sharing partitions hardware to enable concurrent execution. Foundational mechanisms include intra-context CUDA Streams [12] and NVIDIA’s static partitioning tools, MPS and MIG. Systems such as GSlice [15], GPUlets [8], and COLTI [53] build upon these primitives to construct higher-level spatio-temporal abstractions. Yet, because their underlying substrates rely on rigid, static partitioning, they cannot elastically adapt to the highly dynamic resource demands characteristic of modern workloads.
Recent literature pushes for even finer-grained multiplexing. POD-Attention [43] rewrites attention kernels to collocate prefill and decode phases, while LithOS [9] atomizes kernels to pack execution fragments at the Texture Processing Cluster (TPC) level. While effective for utilization, these invasive techniques necessitate user-kernel modifications, risking behavioral inequivalence and complicating offline-to-online correctness guarantees. In contrast, DetShare avoids this semantic trap: by decoupling logical contexts from physical hardware, it dynamically remaps quotas at runtime to achieve elastic, fine-grained spatial multiplexing while strictly preserving execution determinism.
LLM Inference Scheduling.
The unique resource asymmetries of LLM inference have spurred specialized scheduling techniques. Recent work manages the prefill-decode imbalance via stall-free chunking (Sarathi-Serve [1]) and fairness-aware phase balancing (FairBatching [52]). Other efforts explore temporal scheduling through preemptible context migrations (Llumnix [68]) and output-length-guided token-level preemption (TRAIL [63]), or focus on QoS mitigation via adapter-aware caching (Chameleon [37]) and real-time latency attribution (LLMVisor [41], Niyama [24]). While these systems optimize application-level scheduling, DetShare provides a transparent, infrastructure-level foundation. Its extensible policy framework allows operators to plug in diverse algorithms while enforcing execution through strict, hardware-level isolation primitives.
GPU Memory Management and Disaggregation.
Frameworks like AvA [80], rCUDA [17], and DCUDA [28] leverage API remoting for disaggregated execution and flexible GPU kernel placement. Simultaneously, memory-centric systems like DeepUM [42], Zico [48], and TGS [8] optimize device utilization by reducing allocation overheads and shrinking active working-set footprints. These memory optimization and disaggregation techniques are entirely orthogonal and complementary to DetShare, presenting natural avenues for future integration to further maximize efficiency in shared environments.
8 Conclusion
In this paper, we identify and resolve the fundamental challenges of execution determinism, performance predictability, and practical extensibility in modern GPU sharing. To this end, we propose DetShare, a transparent GPU sharing system designed to maximize hardware utilization without compromising determinism. By fundamentally decoupling logical execution contexts from physical hardware resources via a novel GPU coroutine abstraction, DetShare enables flexible spatial multiplexing with zero kernel modifications. Ultimately, DetShare successfully achieves strict semantic and performance determinism, providing a robust, zero-engineering-tax foundation for multi-tenant AI deployments.